# LivePyxel: Accelerating image annotations with a Python-integrated webcam live streaming Uriel Garcilazo-Cruz,a,b,†, Joseph O. Okemea,b, and Rodrigo A. Vargas-Hernándeza,b,c,†

### v1

@@
The lack of flexible annotation tools has hindered the deployment of AI models in some scientific areas. Most existing image annotation software requires users to upload a precollected dataset, which limits support for on-demand pipelines and introduces unnecessary steps to acquire images. This constraint is particularly problematic in laboratory environments, where real-time data acquisition from instruments such as microscopes is increasingly common. In this work, we introduce LivePixel, a Python-based graphical user interface that integrates with imaging systems, such as webcams, microscopes, and others, to enable real-time image annotation. LivePyxel is designed to be easy to use through a simple interface that allows users to precisely delimit areas for annotation using tools commonly found in commercial graphics editing software. Of particular interest is the availability of Bézier splines and binary masks, and the software’s capacity to work with non-destructive layers that enable high-performance editing. LivePyxel also integrates a wide compatibility across video devices, and it’s optimized for object detection operations via the use of OpenCV in combination with high-performance libraries designed to handle matrix and linear algebra operations via Numpy effectively. LivePyxel facilitates seamless data collection and labeling, accelerating the development of AI models in experimental workflows. LivePyxel freely available at https://github.com/UGarCil/LivePyxel
@@ 1 Introduction
The capabilities of vision models (ViM) depend critically on the quality and availability of images1, 2, 3, 4, 5, 6, 7, 6. However, collecting high-quality annotated images is time consuming and this annotation bottleneck 8 negatively impacts the integration of ViM in highly specialized scientific domains, such as cellular imaging 9, 10, civil infrastructure 11, crop profile characterization 12, environmental microscopy 13 and marine conservation 14 to name a few. To address these data limitations, researchers have developed alternative approaches, including data augmentation schemes 15 and more powerful models capable of segmenting and classifying diverse object types with minimal training (e.g., the Segment Anything Model (SAM) 16. However, these techniques still rely on the precision and accuracy of the initial annotations of domain experts. Ensuring high-quality specialized datasets requires the input of these experts 17, which can be both a limiting factor and financially costly 18. The increasing availability of affordable, high-resolution imaging hardware highlights the importance of easy-to-use open-source software. Such tools can lower technical barriers and facilitate broader participation from individuals who possess domain knowledge but lack the expertise to collect and preprocess large datasets systematically.
@@
An image annotation tool focused on usability should also prioritize simple graphical user interface (GUI) components with a shallow learning curve, minimizing the need for technical support or costly training for specialized personnel. Although existing tools like LabelMe 19, VGG Image Annotator (VIA), and COCO Annotator 20 offer pixel-level annotations for segmentation tasks, they lack live camera integration, a critical gap for workflows requiring immediate feedback or iterative labeling. These tools, though flexible in formatting (e.g., COCO JSON, Pascal VOC), often require users to navigate feature-heavy interfaces or offline workflows (e.g., biologists, field technicians). Moreover, most annotation software supports pixel-level annotation only through polygons or rectangular boxes. While effective for rigid geometries, these primitives are poorly suited for organic or curved structures. Approximating a smooth contour with polygons requires a high number of vertices, introducing annotation inefficiency and potential geometric bias that can propagate into downstream vision models. By contrast, graphic design software routinely uses Bézier splines to capture curves with minimal control points and high precision. This approach offers a more natural representation for biological or irregular shapes, making splines a compelling alternative to conventional polygon-based labeling. Commercial platforms such as Labelbox and Supervisely emphasize collaboration features but omit both live annotation and spline support. RectLabel (macOS-only) supports Bézier curves but does not allow real-time input. Web-based tools like CVAT and Label Studio provide scalability yet remain restricted to pre-recorded media. Together, these limitations highlight the absence of an integrated solution that combines live annotation, spline-based precision, and lightweight usability.
@@
We present LivePyxel, developed here, an open-source Python GUI that integrates real-time video device input with Bézier spline-based segmentation, enabling precise pixel-level annotation of curved structures. LivePyxel was initially developed for on-demand annotation of microscopy images, but it also supports any video device accessible via OpenCV and Python. LivePyxel combines a lightweight, accessible interface with flexible annotation tools, including Bézier splines, polygons, and threshold-based masks, making it suitable for segmentation workflows across diverse research domains. The software is freely available at https://github.com/UGarCil/LivePyxel and can be installed through Pip, with installation details found in the latest version of the repository along with tutorials, examples, and additional code.
@@
The paper is structured as follows: Section 2 describes the specification and deployment details of LivePyxel, developed in this work. Section 3 demonstrates its application through an image segmentation task, where annotated images are used to train a ViM 21 in two scenarios: 1) segmenting eight different microorganisms and 2) performing data engineering with binary masks. LivePyxel is designed to streamline data collection and labeling, accelerating the development of AI models in experimental workflows that require real-time data manipulation or large-scale batch processing.
@@ 2 LivePyxel
LivePyxel is a Python-based GUI designed to streamline pixel-level annotation for image segmentation tasks. Its broad compatibility with imaging devices is achieved through the use of the OpenCV 22 library, which enables real-time video input from virtually any camera and streams it directly to the annotation canvas. We chose OpenCV for its efficient handling of images as numerical matrices, leveraging a compiled back-end for performance while maintaining flexibility through seamless integration with NumPy 23 for fast manipulation in Python. The GUI itself is built using the Qt framework, which allows NumPy arrays to be rendered directly as images, resulting in a responsive and user-friendly interface. Fig. 1 is an overview of the LivePyxel architecture.
The paper is structured as follows: Section 2 describes the specification and deployment details of LivePyxel, developed in this work. Section 3 demonstrates its application through an image segmentation task, where annotated images are used to train a ViM 21 in two scenarios: 1) segmenting eight different microorganisms and 2) performing data engineering with binary masks. LivePyxel is designed to streamline data collection and labeling, accelerating the development of AI models in experimental workflows that require real-time data manipulation or large-scale batch processing.
@@
Fig.  1: Overview of the LivePyxel architecture, illustrating the integration of webcam and image inputs with OpenCV for real-time image annotation. The system supports input from live microscopy or camera feeds as well as pre-existing datasets, processes them through the LivePyxel interface, and outputs both annotated images and segmentation masks for storage or further analysis by a vision model.
@@
The LivePyxel graphical user interface (GUI) integrates several interactive components to streamline the annotation workflow. Mask display properties, such as opacity and binary threshold, can be adjusted using the sliders in the control section (Fig. 2-A). Annotation categories are managed in the labels panel (Fig. 2-B), where users can add, edit, or delete classes. The annotation panel (Fig. 2-C) provides high-level controls for switching input sources, capturing frames, and toggling annotation mode. Drawing and editing actions are performed using the toolbar (Fig. 2-D), which offers tools for creating, modifying, or erasing masks. The canvas (Fig. 2-E) displays either live microscope or uploaded images, over which masks are layered non-destructively in real time. Navigation buttons (Fig. 2-F) allow users to cycle through frames or dataset entries. Once annotations are complete, LivePyxel saves both the annotated images and corresponding masks into automatically generated subfolders within a user-specified directory. In addition to real-time webcam annotation, LivePyxel also supports traditional workflows by allowing users to upload pre-existing image datasets that follow the required folder structure.
@@
Fig.  2: Preview of the LivePyxel graphical user interface (GUI) with its main components labeled: (A) sliders for adjusting mask opacity and binary mask threshold, (B) labels panel for managing annotation categories, (C) annotation panel with main control buttons, (D) toolbar for selecting drawing and editing tools, (E) central canvas where objects are annotated on microscopy images in real time, and (F) navigation controls for browsing through image frames or dataset entries.
@@ 2.1 Bézier splines for non-tradictional curvatures
A central feature of LivePyxel is its support for Bézier splines 24; a mathematical representation of curves that originated in computer graphics and has become an industry standard for precision editing. Unlike traditional polygon tools, which approximate shapes using straight-line segments and are better suited for rigid structures like crystals, buildings, or mechanical parts, Bézier splines allow for smooth, continuous contours; see Fig. 3. This makes them ideal for segmenting organic, curved shapes typically encountered in biological datasets, such as cells, tissues, or protozoa.
Each Bézier unit is defined by three control points: two end points and a central handle that determines curvature; see Fig. 3. By connecting multiple units, users can construct complex outlines with high fidelity to natural forms. This modular structure offers fine control over both curvature and sharp transitions. The advantage of splines for delimiting contours is illustrated by SplineDist 25, which extends over the popular StarDist framework by modeling objects as planar parametric spline curves, allowing more flexible and smooth segmentation boundaries and solving issues with non-convex geometries. Moreover, the use of splines in the preparation of annotated data for segmentation tasks has been documented in clinical applications 26, 27. In biomedical imaging, the use of splines enables users to trace smooth cell membranes with highly organic contours and tightly coiled or angular biological features, something that would require excessive effort and precision with polygon tools. The ability to produce accurate, pixel-level masks with fewer interactions not only improves annotation speed but also reduces user fatigue and annotation bias. As a result, Bézier splines in LivePyxel improve both the efficiency and the quality of segmentation in tasks where precision is critical, such as microscopy, radiography, and digital morphology.
Fig.  3: Comparison between polygon (panels A and C) and Bézier-splines (panels B and D) in the delimitation of boundaries of a tardigrade’s body. One Bézier unit is defined by three points. ptA and ptB constitute the start and end of the unit, whereas ptG serves as a point of gravity to pull the line away from the user’s cursor, forcing it to become a curve.
@@ 2.2 Binary Masks and Multi-Layered Structure
Another important feature provided by LivePyxel is the generation and manipulation of binary masks, which enable semi-automated annotation in highly controlled imaging environments. LivePyxel generates binary masks by applying pixel-level operations using OpenCV throughout the frame, then assigning a value of 0 or $1$ to every pixel before creating an annotation mask. Using this technique, the user can isolate regions of high contrast against a uniform background, such as a white or black field. This proves to be especially useful in biomedical applications involving solid and organic objects 28, 29, and potentially significant in many other fields of biological science, where annotation targets are organic and contour-heavy. We showcase this feature in Section 3.2.
@@
Fig.  4: Illustration of the multi-layer annotation process. Each annotation is stored as a separate layer within a stack, which is subsequently merged into a single raster image from bottom to top upon saving.
@@ 2.3 Webcam integration
LivePyxel, through the OpenCV library, integrates the webcam feed into the GUI. The program begins by searching for up to four camera devices connected to the computer, displaying the first device found. This multicamera feature is useful in situations where the user has multiple video-device stations connected to the application and may require flexibility in accessing different cameras in controlled environments. Once loaded, webcam images are rendered onto the canvas, allowing users to perform annotations either in real-time during live streaming or on captured frames. The user concludes an annotation by clicking on the Annotate button. The annotation will be stored in the folder specified by the user, in the form of an RGB image in the native webcam’s resolution, and a matching mask with the pixel annotations encoded by the colors chosen by the user. LivePyxel is capable of reading most formats supported by OpenCV, including .jpg, .png, .tiff, and .bmp, and the annotated images are saved in .png format. The identity and color of each labeled class are stored in a .json file for its readability and its common use in a wide range of programming and app development applications.
@@ 2.4 Tablet integration
The use of drawing tablets in combination with graphics editors is a desirable practice as it provides notable advantages over the use of a computer mouse. Tablet integration improves annotations by reducing the time required to delimit the contour of objects without compromising precision. The libraries that integrate the core functionality of the program, OpenCV and pyQt, contain highly optimized compatibility with computer mouse-like devices, resulting in a smooth “plug-and-play" integration with LivePyxel (Fig. 5). The “Free Hand" is a specific tool designed to improve the user experience with drawing tablets. It allows the user to automatically generate new vertices in a polygon by calculating the distance between the last vertex and the position of the tablet’s stylus, adding a new vertex when the distance is greater than a reference threshold. This behavior allows the user to focus entirely on the hand movement, tracking the contour. This functionality was essential for creating the dataset used to train a vision model to segment eight different microorganisms in Section 3.1, using an Artist Display 15.6 Pro tablet.
Fig.  5: The free annotation tool enables users to automatically trace the contour of an annotation target by delegating vertex placement to the program, which positions them based on distance thresholds.
@@ 3 Examples
Here, we present two distinct examples demonstrating the use of LivePyxel for labeling training data in pixel-level vision tasks. In Section 3.1, we showcase its application in a segmentation task involving eight different microorganisms, highlighting manual data annotation using Bézier splines and other built-in tools. In Section 3.2, we describe the development of a large dataset through semi-automated labeling with binary masks, followed by data augmentation techniques to evaluate LivePyxel’s scalability in dataset creation for training vision models. For both examples, the training pipelines and models were implemented using PyTorch 30.

### v2

@@
The lack of flexible annotation tools has hindered the deployment of AI models in some scientific areas. Most existing image annotation software requires users to upload a precollected dataset, which limits support for on-demand pipelines and introduces unnecessary steps to acquire images. This constraint is particularly problematic in laboratory environments, where on-site data acquisition from instruments such as microscopes is increasingly common. In this work, we introduce LivePixel, a Python-based graphical user interface that integrates with imaging systems, such as webcams, microscopes, and others, to enable on-site image annotation. LivePyxel is designed to be easy to use through a simple interface that allows users to precisely delimit areas for annotation using tools commonly found in commercial graphics editing software. Of particular interest is the availability of Bézier splines and binary masks, and the software’s capacity to work with non-destructive layers that enable high-performance editing. LivePyxel also integrates a wide compatibility across video devices, and it’s optimized for object detection operations via the use of OpenCV in combination with high-performance libraries designed to handle matrix and linear algebra operations via Numpy effectively. LivePyxel facilitates seamless data collection and labeling, accelerating the development of AI models in experimental workflows. LivePyxel is freely available at https://github.com/UGarCil/LivePyxel
footnotetext: a Department of Chemistry and Chemical Biology, McMaster University, Hamilton, ON, Canada
b School of Computational Science and Engineering, McMaster University, Hamilton, ON, Canada
footnotetext: c Brockhouse Institute for Materials Research, McMaster University, Hamilton, ON, Canada
† Corresponding author: garcilau@mcmaster.ca, vargashr@mcmaster.ca
@@ 1 Introduction
The capabilities of vision models (ViM) depend critically on the quality and availability of imagesBilalEtAl2024, chi2020deep, KARIMI2020101759, singh2024shifting, rs15204987, taran2020impact, djuravs2024dataset, taran2020impact. However, collecting high-quality annotated images is time consuming and this annotation bottleneck litjens2017survey negatively impacts the integration of ViM in highly specialized scientific domains, such as cellular imaging tajbakhsh2020embracing, moen2019deep, civil infrastructure Yuan2024Review, crop profile characterization agronomy15051157, environmental microscopy Luo_et_al_2018 and marine conservation sauder2025coralscapes to name a few. To address these data limitations, researchers have developed alternative approaches, including data augmentation schemes wang2021annotation and more powerful models capable of segmenting and classifying diverse object types with minimal training (e.g., the Segment Anything Model (SAM) kirillov2023segany. However, these techniques still rely on the precision and accuracy of the initial annotations of domain experts. Ensuring high-quality specialized datasets requires the input of these experts Nodine1999, which can be both a limiting factor and financially costly hosny2018artificial. The increasing availability of affordable, high-resolution imaging hardware highlights the importance of easy-to-use open-source software. Such tools can lower technical barriers and facilitate broader participation from individuals who possess domain knowledge but lack the expertise to collect and preprocess large datasets systematically.
@@
An image annotation tool focused on usability should also prioritize simple graphical user interface (GUI) components with a shallow learning curve, minimizing the need for technical support or costly training for specialized personnel. Although existing tools like LabelMe russell2008labelme, VGG Image Annotator (VIA) dutta2019vgg, and COCO Annotator cocoannotator offer pixel-level annotations for segmentation tasks, they lack live camera integration, a critical gap for workflows requiring immediate feedback or iterative labeling. These tools, though flexible in formatting (e.g., COCO JSON, Pascal VOC), often require users to navigate feature-heavy interfaces or offline workflows (e.g., biologists, field technicians). Moreover, most annotation software supports pixel-level annotation only through polygons or rectangular boxes. While effective for rigid geometries, these primitives are poorly suited for organic or curved structures. Approximating a smooth contour with polygons requires a high number of vertices, introducing annotation inefficiency and potential geometric bias that can propagate into downstream vision models. By contrast, graphic design software routinely uses Bézier splines to capture curves with minimal control points and high precision. This approach offers a more natural representation for biological or irregular shapes, making splines a compelling alternative to conventional polygon-based labeling. Commercial platforms such as Labelbox and Supervisely emphasize collaboration features but omit both live annotation and spline support. RectLabel (macOS-only) supports Bézier curves but does not allow on-site input. Web-based tools like CVAT cvat_2023 and Label Studio labelstudio_2020 provide scalability yet remain restricted to pre-recorded media. Together, these limitations highlight the absence of an integrated solution that combines live annotation, spline-based precision, and lightweight usability.
@@
We present LivePyxel, developed here, an open-source Python GUI that integrates on-site video device input with Bézier spline-based segmentation, enabling precise pixel-level annotation of curved structures. LivePyxel was initially developed for on-demand annotation of microscopy images, but it also supports any video device accessible via OpenCV and Python. LivePyxel combines a lightweight, accessible interface with flexible annotation tools, including Bézier splines, polygons, and threshold-based masks, making it suitable for segmentation workflows across diverse research domains. The software is freely available at https://github.com/UGarCil/LivePyxel and can be installed through PyPI, with installation details found in the latest version of the repository along with tutorials, examples, and additional code.
@@
The paper is structured as follows: Section 2 describes the specification and deployment details of LivePyxel, developed in this work. Section 3 demonstrates its application through an image segmentation task, where annotated images are used to train a ViM ronneberger2015u in two scenarios: 1) segmenting eight different microorganisms and 2) performing data engineering with binary masks. LivePyxel is designed to streamline data collection and labeling, accelerating the development of AI models in experimental workflows that require on-site data manipulation or large-scale batch processing.
@@ 2 LivePyxel
LivePyxel is composed of an ecosystem of Python scripts that work cohesively to provide users with a graphical interface. This architecture provides a tool with a higher level of abstraction compared to most data science workflows, typically executed via Jupyter Notebooks. The abstraction level is higher because the code that makes the software does not directly encode the annotation process, but rather builds interoperability across objects that make up the graphical user interface (GUI). This architecture was chosen because the domain of the task addressed by the program naturally belongs to a graphics editor, and because graphical interfaces make the tool available to a wider audience of users with very little experience in programming and code.
@@
The broad compatibility of LivePyxel with imaging devices is achieved through the use of the OpenCV bradski2000opencv library, which enables on-site video input from virtually any camera and streams it directly to the annotation canvas. We chose OpenCV for its efficient handling of images as numerical matrices, leveraging a compiled back-end for performance while maintaining flexibility through seamless integration with NumPy harris2020array for fast manipulation in Python. The GUI itself is built using the Qt framework, which allows NumPy arrays to be rendered directly as images, resulting in a responsive and user-friendly interface. Fig. 1 is an overview of the LivePyxel architecture.
@@
Fig. 1: Overview of the LivePyxel architecture, illustrating the integration of webcam and image inputs with OpenCV for on-site image annotation. The system supports input from live microscopy or camera feeds as well as pre-existing datasets, processes them through the LivePyxel interface, and outputs both annotated images and segmentation masks for storage or further analysis by a vision model.
@@
The LivePyxel GUI integrates several interactive components to streamline the annotation workflow. Mask display properties, such as opacity and binary threshold, can be adjusted using the sliders in the control section (Fig. 2-A). Annotation categories are managed in the labels panel (Fig. 2-B), where users can add, edit, or delete classes. The annotation panel (Fig. 2-C) provides high-level controls for switching input sources, capturing frames, and toggling annotation mode. Drawing and editing actions are performed using the toolbar (Fig. 2-D), which offers tools for creating, modifying, or erasing masks. The canvas (Fig. 2-E) displays either live microscope or uploaded images, over which masks are layered non-destructively in real time. Navigation buttons (Fig. 2-F) allow users to cycle through frames or dataset entries. Once annotations are complete, LivePyxel saves both the annotated images and corresponding masks into automatically generated subfolders within a user-specified directory. In addition to on-site webcam annotation, LivePyxel also supports traditional workflows by allowing users to upload pre-existing image datasets that follow the required folder structure.
@@
Fig. 2: Preview of the LivePyxel graphical user interface (GUI) with its main components labeled: (A) sliders for adjusting mask opacity and binary mask threshold, (B) labels panel for managing annotation categories, (C) annotation panel with main control buttons, (D) toolbar for selecting drawing and editing tools, (E) central canvas where objects are annotated on microscopy images in real time, and (F) navigation controls for browsing through image frames or dataset entries.
@@ 2.1 Bézier splines for non-tradictional curvatures
A central feature of LivePyxel is its support for Bézier splines rogers2000introduction; a mathematical representation of curves that originated in computer graphics and has become an industry standard for precision editing. Unlike traditional polygon tools, which approximate shapes using straight-line segments and are better suited for rigid structures like crystals, buildings, or mechanical parts, Bézier splines allow for smooth, continuous contours; see Fig. 3. This makes them ideal for segmenting organic, curved shapes typically encountered in biological datasets, such as cells, tissues, or protozoa.
Each Bézier unit is defined by three control points: two end points and a central handle that determines curvature; see Fig. 3-B. By connecting multiple units, users can construct complex outlines with high fidelity to natural forms. This modular structure offers fine control over both curvature and sharp transitions. The advantage of splines for delimiting contours is illustrated by SplineDist mandal2021splinedist, which extends over the popular StarDist framework by modeling objects as planar parametric spline curves, allowing more flexible and smooth segmentation boundaries and solving issues with non-convex geometries. Moreover, the use of splines in the preparation of annotated data for segmentation tasks has been documented in clinical applications huellebrand2022collaborative, materka2024using. In biomedical imaging, the use of splines enables users to trace smooth cell membranes with highly organic contours and tightly coiled or angular biological features, something that would require excessive effort and precision with polygon tools. The ability to produce accurate, pixel-level masks with fewer interactions not only improves annotation speed but also reduces user fatigue and annotation bias. As a result, Bézier splines in LivePyxel improve both the efficiency and the quality of segmentation in tasks where precision is critical, such as microscopy, radiography, and digital morphology.
Fig. 3: Comparison between polygon (panels A and C) and Bézier-splines (panels B and D) in the delimitation of boundaries of a tardigrade’s body. One Bézier unit is defined by three points. ptA and ptB constitute the start and end of the unit, whereas ptG serves as a point of gravity to pull the line away from the user’s cursor, forcing it to become a curve.
@@ 2.2 Binary Masks and Multi-Layered Structure
Another important feature provided by LivePyxel is the generation and manipulation of binary masks, which enable semi-automated annotation in highly controlled imaging environments. LivePyxel generates binary masks by applying pixel-level operations using OpenCV throughout the frame, then assigning a value of 0 or $1$ to every pixel before creating an annotation mask. Using this technique, the user can isolate regions of high contrast against a uniform background, such as a white or black field. This proves to be especially useful in biomedical applications involving solid and organic objects chen2023bezierseg, galvez2021nurbs, and potentially significant in many other fields of biological science, where annotation targets are organic and contour-heavy. We showcase this feature in Section 3.2.
@@
Fig. 4: Illustration of the multi-layer annotation process. Each annotation is stored as a separate layer within a stack, which is subsequently merged into a single raster image from bottom to top upon saving.
@@ 2.3 Webcam integration
LivePyxel, through the OpenCV library, integrates the webcam feed into the GUI. The program begins by searching for up to four camera devices connected to the computer, displaying the first device found. This multicamera feature is useful in situations where the user has multiple video-device stations connected to the application and may require flexibility in accessing different cameras in controlled environments. Once loaded, webcam images are rendered onto the canvas, allowing users to perform annotations either in on-site during live streaming or on captured frames. The user concludes an annotation by clicking on the Annotate button. The annotation will be stored in the folder specified by the user, in the form of an RGB image in the native webcam’s resolution, and a matching mask with the pixel annotations encoded by the colors chosen by the user. LivePyxel is capable of reading most formats supported by OpenCV, including .jpg, .png, .tiff, and .bmp, and the annotated images are saved in .png format. The identity and color of each labeled class are stored in a .json file for its readability and its common use in a wide range of programming and app development applications.
@@ 2.4 Tablet integration
The use of drawing tablets in combination with graphics editors is a desirable practice as it provides notable advantages over the use of a computer mouse. Tablet integration improves annotations by reducing the time required to delimit the contour of objects without compromising precision. The libraries that integrate the core functionality of the program, OpenCV and pyQt, contain highly optimized compatibility with computer mouse-like devices, resulting in a smooth “plug-and-play" integration with LivePyxel. The “Free Hand" is a specific tool designed to improve the user experience with drawing tablets. It allows the user to automatically generate new vertices in a polygon by calculating the distance between the last vertex and the position of the tablet’s stylus, adding a new vertex when the distance is greater than a reference threshold. This behavior allows the user to focus entirely on the hand movement, tracking the contour. This functionality was essential for creating the dataset used to train a vision model to segment eight different microorganisms in Section 3.1, using an Artist Display 15.6 Pro tablet.
@@ 3 Examples
Here, we present two distinct examples demonstrating the use of LivePyxel for labeling training data in pixel-level vision tasks. In Section 3.1, we showcase its application in a segmentation task involving eight different microorganisms, highlighting manual data annotation using Bézier splines and other built-in tools. In Section 3.2, we describe the development of a large dataset through semi-automated labeling with binary masks, followed by data augmentation techniques to evaluate LivePyxel’s scalability in dataset creation for training vision models. A video introduction to the use of binary masks for the semi-automation of dataset annotations is presented at: tutorial link. For both examples, the training pipelines and models were implemented using PyTorch pytorch. Finally, in Section 3.3, we compared the labeling accuracy of LivePixel with four other annotation tools.